A sample of a compound consists of C , H and N only.If 24 gm of carbon , 6.022 x 1024 atoms of hydrogen and 2 -gm atom of nitrogen is present and vapour density of the compound is 16.5 , then its molecular formula will be ?

Grade:

7 Answers

Shivam Dimri
11 years ago

vapor density is 16.5

=> molecular mass should be 16.5 * 2 = 33

Now it has 24 gm carbon => 2x moles of carbon

2-gm atom of N = 2x moles of nitrogen

.

.

.

going this way the empirical mass> molecular mass

that immpossible

you posted a wrong question

check again!!

 

Akash Kumar Dutta
11 years ago

given vapour density=16.5
so molecular mass= 2.d=33 gm
no. of moles of carbon=2
no. of moles of hydrogen=10
moles of nitrogen=2,
ratio of moles=2:10:2=1:5:1
but according to this the mass comes=31 gm.
hence the no of hydrogen req. more=2
so the formula becomes....
            CH(7)N

tejas naresh patel
11 years ago

the method i try to remove empirical formula

shiddhant bhattacharya
11 years ago

is it CH5N

Saurabh Anand
11 years ago

vapour density = molecular mass/2

so ur ans is....16.5*2 = 33g/mol

FITJEE
10 years ago

CH5N

vishal ojha
10 years ago

wrong question